Find ∆H in terms of kj for the following reaction
M(g)+2X(g)--->M²+(g)+2X^-(g)
(IE)1 of M(g)=705Kj/mol
(IE)2 of M(g)=1151KJ/mol
(EA)1 of X(g)= -328KJ/mol
Answer 1200KJ
1
Expert's answer
2019-08-09T05:17:41-0400

∆H=((IE)1(M) +(IE)2(M)+2(EA)1(X))-(2·∆H(X)+∆H(M))=705+1151+2·(-328)-2·0+0=1200 KJ